# Offline Clustering of Linear Bandits: Unlocking the Power of Clusters in Data-Limited Environments

### v1

@@ D.1 Detailed Algorithm: Off-CLUB
1:   Input: User $u_{\text{test}}$, action set $A_{\text{test}}$, dataset $\mathcal{D}$, parameters $\alpha$, $\lambda>1$ and $\delta>0$
2:   Initialization: Construct a complete graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, and compute $M_{u}$, $\bm{b}_{u}$, $\hat{\bm{\theta}}_{u}$, and $\text{CI}_{u}$ as in (1) for each user $u\in\mathcal{V}$
@@
7:  Let $\tilde{\mathcal{G}}=(\mathcal{V},\tilde{\mathcal{E}})$ denote the updated graph
@@
$\bm{a}_{\text{test}}=\mathop{\mathrm{argmax}}\limits_{\bm{a}\in A_{\text{test}}}\left(\left(\tilde{\bm{\theta}}_{u_{\text{test}}}\right)^{\top}\bm{a}-\beta\left\|\bm{a}\right\|_{\left(\tilde{M}_{u_{\text{test}}}\right)^{-1}}\right)$
@@
The algorithm takes similar inputs to Algorithm 1, except that it does not require $\hat{\gamma}$. It starts with a complete graph $\mathcal{G}$ and iterates through all edges $(u_{1},u_{2})\in\mathcal{E}$ during the Cluster Phase. If the condition $\left\|\hat{\bm{\theta}}_{u_{1}}-\hat{\bm{\theta}}_{u_{2}}\right\|_{2}>\alpha\left(\text{CI}_{u_{1}}+\text{CI}_{u_{2}}\right)$ holds, the algorithm concludes that users $u_{1}$ and $u_{2}$ do not share the same preference vector and removes the edge between them (Line 5). After processing all edges, the resulting graph $\tilde{\mathcal{G}}$ separates users with different preference vectors into distinct clusters, provided that each user has sufficient samples to ensure a low confidence interval. However, if a user $u$ has fewer samples in $\mathcal{D}_{u}$, the resulting higher confidence interval $\text{CI}_{u}$ reduces the likelihood of satisfying condition (7), potentially leading to the inclusion of heterogeneous users in the same cluster. The conditions necessary to ensure the algorithm’s correctness are discussed in detail in Section 5.2.
@@
Similar to Algorithm 1, this algorithm aggregates data from each user and its neighbors in $\tilde{\mathcal{G}}$ to form a new dataset and computes the necessary statistics (Lines 9 and 10). It is important to note that this data aggregation strategy, like that in Algorithm 1, differs from the online version of the CLUB algorithm introduced in [Gentile et al., 2014], where data from all users within the same connected component of the graph is utilized. For instance, consider a scenario where $(u,v)\in\tilde{\mathcal{E}}$ and $(v,w)\in\tilde{\mathcal{E}}$, but $(u,w)\notin\tilde{\mathcal{E}}$. In this case, samples from $\mathcal{D}_{w}$ are excluded from $\tilde{\mathcal{D}}_{u}$ in Algorithm 2, whereas the online version includes them in $\tilde{\mathcal{D}}_{u}$. The restricted aggregation in Algorithm 2 is crucial in the offline setting: since users that are no longer neighbors of $u$ are guaranteed not to share the same preference vector, including their data in $\tilde{\mathcal{D}}_{u}$ would introduce bias. In the online setting, however, such bias is not a concern because infinite online data ensures that all edges between users with different preference vectors are reliably removed. This distinction highlights the importance of careful data selection in the offline setting to maintain prediction accuracy.
Finally, the Decision Phase is performed in a manner analogous to Algorithm 2, using a pessimistic estimate to select $\bm{a}_{\text{test}}$.
D.2 Additional Remarks
Remark D.1 (Asymptotic Properties).
@@ D.2 Additional Remarks
In the special case where $\mathcal{V}(j_{u_{\textnormal{test}}})=\{u_{\textnormal{test}}\}$ (i.e., the test user is distinct from all other users in $\mathcal{V}$), the result reduces to the standard bound for linear bandits, $\tilde{O}\left(\sqrt{\frac{d}{N_{u_{\textnormal{test}}}}}\right)$. Conversely, when $\mathcal{V}(j_{u_{\textnormal{test}}})\approx\mathcal{V}$ (i.e., most users share the same preference vector), Algorithm 2 effectively aggregates data from nearly all users, leveraging the large sample size to make more accurate predictions.
Remark D.2 (Intuition of Assumption 5.1).
@@
The key assumption underlying Theorem 5.2 is data sufficiency. As defined in Assumption 5.1, complete data sufficiency comprises two main components. The first component depends on $1/\lambda_{a}^{2}$, which ensures sufficient information about each dimension of the preference vector for each user. Without this level of data, the algorithm cannot reliably estimate each dimension, making accurate predictions infeasible. In most cases, $\lambda_{a}$ can be treated as a constant, rendering the first term effectively constant. The second component in Assumption 5.1 depends on $1/\gamma^{2}$, which is the main part of the data sufficiency requirement. This term guarantees that the true cluster of each user can be accurately identified. Insufficient data in this regard prevents the algorithm from correctly assigning users to their respective clusters, potentially leading to the inclusion of users with different preference vectors and introducing bias into the predictions. Notably, a smaller $\gamma$ corresponds to a stricter data sufficiency requirement.

### v2

@@ F.1 Detailed Algorithm: Off-CLUB
1:Input: User $u_{\text{test}}$, action set $\mathcal{A}_{\text{test}}$, dataset $\mathcal{D}$, parameters $\alpha$, $\lambda>1$ and $\delta>0$
2:Initialization: Construct a complete graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where $\mathcal{V}=\mathcal{U}$ is the set of all users, and compute $M_{u}$, $\bm{b}_{u}$, $\hat{\bm{\theta}}_{u}$, and $\text{CI}_{u}$ as in (1) for each user $u\in\mathcal{V}$
@@
[…]
@@
$\bm{a}_{\text{test}}=\mathop{\mathrm{argmax}}\limits_{\bm{a}\in\mathcal{A}_{\text{test}}}\left(\left(\tilde{\bm{\theta}}_{u_{\text{test}}}\right)^{\top}\bm{a}-\beta\left\|\bm{a}\right\|_{\left(\tilde{M}_{u_{\text{test}}}\right)^{-1}}\right)$
@@
Similar to Algorithm 1, this algorithm aggregates data from each user and its neighbors in $\tilde{\mathcal{G}}$ to form a new dataset and computes the necessary statistics (Lines 9 and 10). It is important to note that this data aggregation strategy, like that in Algorithm 1, differs from the online version of the CLUB algorithm introduced in Gentile et al. [2014], where data from all users within the same connected component of the graph is utilized. For instance, consider a scenario where $(u,v)\in\tilde{\mathcal{E}}$ and $(v,w)\in\tilde{\mathcal{E}}$, but $(u,w)\notin\tilde{\mathcal{E}}$. In this case, samples from $\mathcal{D}_{w}$ are excluded from $\tilde{\mathcal{D}}_{u}$ in Algorithm 2, whereas the online version includes them in $\tilde{\mathcal{D}}_{u}$. The restricted aggregation in Algorithm 2 is crucial in the offline setting: since users that are no longer neighbors of $u$ are guaranteed not to share the same preference vector, including their data in $\tilde{\mathcal{D}}_{u}$ would introduce bias. In the online setting, however, such bias is not a concern because infinite online data ensures that all edges between users with different preference vectors are reliably removed. This distinction highlights the importance of careful data selection in the offline setting to maintain prediction accuracy.
Finally, the Decision Phase is performed in a manner analogous to Algorithm 2, using a pessimistic estimate to select $\bm{a}_{\text{test}}$. This pessimistic strategy differs from the optimistic approaches used in prior online clustering of bandits [Gentile et al., 2014, Li and Zhang, 2018, Li et al., 2019, Wang et al., 2025, Li et al., 2025], which balance the exploration–exploitation tradeoff. Instead, it follows the pessimism principle widely adopted in offline bandits and RL [Jin et al., 2021, Li et al., 2022, Liu et al., 2025], avoiding over-reliance on poorly represented dimensions.
F.2 Additional Remarks
Remark F.1 (Asymptotic Properties).
@@ F.2 Additional Remarks
In the special case where $\mathcal{V}(j_{u_{\textnormal{test}}})=\{u_{\textnormal{test}}\}$ (i.e., the test user is distinct from all other users in $\mathcal{V}$), the result reduces to the standard bound for linear bandits, $\tilde{O}\left(\sqrt{\frac{d}{\tilde{\lambda}_{a}N_{u_{\textnormal{test}}}}}\right)$. Conversely, when $\mathcal{V}(j_{u_{\textnormal{test}}})\approx\mathcal{V}$ (i.e., most users share the same preference vector), Algorithm 2 effectively aggregates data from nearly all users, leveraging the large sample size to make more accurate predictions.
Remark F.2 (Intuition of Assumption 5.1).
@@
The key assumption underlying Theorem 5.2 is data sufficiency. As defined in Assumption 5.1, complete data sufficiency comprises two main components. The first component depends on $1/\tilde{\lambda}_{a}^{2}$, which ensures sufficient information about each dimension of the preference vector for each user. Without this level of data, the algorithm cannot reliably estimate each dimension, making accurate predictions infeasible. The second component in Assumption 5.1 depends on $1/\gamma^{2}$, which is the main part of the data sufficiency requirement. This term guarantees that the true cluster of each user can be accurately identified. Insufficient data in this regard prevents the algorithm from correctly assigning users to their respective clusters, potentially leading to the inclusion of users with different preference vectors and introducing bias into the predictions. Notably, a smaller $\gamma$ corresponds to a stricter data sufficiency requirement.